write down degeneracy for isotropic 1 d harmonic oscillator.
1
Expert's answer
2020-11-27T00:41:02-0500

The degeneracy for an p-dimensional quantum harmonic oscillator is given as


"g(n,p)=\frac{(n+p-1)!}{n!(p-1)!}"

it corresponds to the number of weak compositions of the integer n into p integers.

The degeneracy for isotropic 1 d harmonic oscillator is


"g(n)=\frac{(n+1-1)!}{n!(1-1)!}=\frac{(n)!}{n!(0)!}=1"